# A Universal Analytic Model for Gravitational Lensing by Self-Interacting Dark Matter Halos

### v1

@@ 6 Impact of baryons
Figure 13: The distribution of gravothermal phases, represented by color-coded circles with sizes proportional to halo masses, on the size-density plane ($\rho_{H}/\rho_{s}-r_{H}/r_{s}$) of baryons relative to dark matter. The SIDM model considered here assumes $\sigma_{0}/m=100\leavevmode\nobreak\ \rm cm^{2}/g$ and $w=60\leavevmode\nobreak\ \rm km/s$, parameterizing a Rutherford-like scattering cross section. The left and right panels show results for halos excluding and including baryon effects, respectively, highlighting the significant boost in the gravothermal phase due to the presence of baryons. The halo sample is randomly selected from isolated halos with masses exceeding $10^{10}\rm M_{\odot}$ in the TNG-50-1 simulation, with the most massive halo reaching approximately $10^{13}\rm M_{\odot}$.
@@
where $\sigma_{0}/m=100\leavevmode\nobreak\ \rm cm^{2}/g$ and $w=60\leavevmode\nobreak\ \rm km/s$. Strictly speaking, this model applies to distinguishable initial states. See ref. [64] for a consistent particle interpretation in N-body simulations. The results, plotted on the size-density plane ($\rho_{H}/\rho_{s}-r_{H}/r_{s}$), demonstrate that baryons can significantly enhance $\tau$ values, particularly for halos with a high baryon-to-halo mass ratio, which spread towards the upper-right region of the plane.
@@ 7 Conclusion and discussion
To assess the accuracy of our model, we adopt a perspective that separates uncertainties into two categories: those arising from the parametrized hypersurface and those associated with the interfacing parameters. Our calibrated hypersurface itself is highly accurate and representative: for any SIDM halo composed solely of dark matter, we expect it to provide a reliable fit. On the other hand, uncertainty lies in the interfacing parameters—$\rho_{s}$, $r_{s}$, and $\tau$—can be much larger. Specifically, given an NFW halo with fixed $\rho_{s}$ and $r_{s}$, the determination of $\tau$ depends on estimating the core-collapse time (Eq. 2.1), which introduces significant uncertainty. In the literature, this uncertainty is typically parameterized by a normalization factor $C$, which has approximately 10–30% uncertainty. Accurately determining the gravothermal phase remains an open challenge warranting further study.
@@
Beyond theoretical applications, our model is well-suited for studying halos with minimal or negligible baryonic content. Observationally, dark halos can still leave detectable imprints, such as distortions in strong lensing images of host halos, gravitational microlensing effects, and perturbations in stellar streams. Investigating these signatures in the context of SIDM could provide crucial insights into the interacting nature of dark matter. Continued efforts in this direction are essential to unveiling these hidden properties.
@@ Acknowledgments
D.Y. was supported in part by the National Key Research and Development Program of China (No. 2022YFF0503304), the Project for Young Scientists in Basic Research of the Chinese Academy of Sciences (No. YSBR-092). N.L. acknowledges the support of the science research grants from the China Manned Space Project (No. CMS-CSST-2021-A01), the Ministry of Science and Technology of China (No. 2020SKA0110100), and, the CAS Project for Young Scientists in Basic Research (No. YSBR-062). G.L. acknowledges the support from the Natural Science Foundation of China (NSFC) (No. 12333001).

### v2

@@ 6 Impact of baryons
Figure 13: The distribution of gravothermal phases, represented by color-coded circles with sizes proportional to halo masses, on the size-density plane ($\rho_{H}/\rho_{s}-r_{H}/r_{s}$) of baryons relative to dark matter. The SIDM model considered here assumes $\sigma_{0}/m=100\leavevmode\nobreak\ \rm cm^{2}/g$ and $w=60\leavevmode\nobreak\ \rm km/s$, parameterizing a Rutherford-like scattering cross section. The left and right panels show results for halos excluding and including baryon effects, respectively, highlighting the significant boost in the gravothermal phase due to the presence of baryons. The halo sample is randomly selected from isolated halos with masses exceeding $10^{10}\rm M_{\odot}$ in the TNG-50-1 simulation, with the most massive halo reaching approximately $10^{13}\rm M_{\odot}$. With baryonic acceleration included, the fraction of halos that reach the core-collapsed phase ($\tau\geq 1.0$) increases from 1.2% to 21.2%.
@@
where $\sigma_{0}/m=100\leavevmode\nobreak\ \rm cm^{2}/g$ and $w=60\leavevmode\nobreak\ \rm km/s$. Strictly speaking, this model applies to distinguishable initial states. See ref. [64] for a consistent particle interpretation in N-body simulations. The results, plotted on the size-density plane ($\rho_{H}/\rho_{s}-r_{H}/r_{s}$), demonstrate that baryons can significantly enhance $\tau$ values, particularly for halos with a high baryon-to-halo mass ratio, which spread towards the upper-right region of the plane. Overall, the fraction of halos that reach the core-collapsed phase ($\tau\geq 1.0$) increases significantly due to baryonic effects, rising from 1.2% to 21.2%. This enhancement is most prominent in intermediate-mass halos and becomes less pronounced toward both lower-mass dwarf halos and higher-mass cluster halos—represented by smaller and larger blue circles, respectively—where the impact of baryons diminishes. In these mass regimes, our analytic lensing model remains more reliable. Quantitatively, within the mass range $10^{10}-10^{11}\rm M_{\odot}$, the fraction of core-collapsed halos increases from 1.4% in the absence of baryons to 19% when baryonic effects are included. For higher mass bins, $(10^{11},10^{12})\rm M_{\odot}$ and $(10^{12},10^{13})\leavevmode\nobreak\ \rm M_{\odot}$, no core-collapsed halos are found without baryons, but the inclusion of baryons raises the fractions to 41% and 5.0%, respectively.
@@ 7 Conclusion and discussion
[…]
@@
The application of our model goes beyond SIDM. Any halo whose density profile can be fitted by the PSIDM model at a given $\tau$ (Eqs. 3.9, 3.10) can use our framework to obtain analytic lensing predictions. For example, Ref. [92] recently showed that two-component dark matter models with mass segregation can yield halo profiles well approximated by PSIDM. It is also possible to superimpose multiple PSIDM profiles to model more complex halo profiles. The lensing effects can then be obtained by summing the contributions from the individual components.
Beyond theoretical applications, our model is well-suited for studying halos with minimal or negligible baryonic content. Observationally, dark halos can still leave detectable imprints through lensing, including flux ratio anomalies, distortions in the strong lensing images of host galaxies, and gravitational microlensing effects [93, 10, 11, 12, 94, 95, 32].Investigating these signatures in the context of SIDM could provide crucial insights into the interacting nature of dark matter. Continued efforts in this direction are essential to unveiling these hidden properties.
@@ Acknowledgments
D.Y. was supported in part by the National Key Research and Development Program of China (No. 2022YFF0503304), the Project for Young Scientists in Basic Research of the Chinese Academy of Sciences (No. YSBR-092). N.L. acknowledges the support of the science research grants from the China Manned Space Project (No. CMS-CSST-2021-A01), the Ministry of Science and Technology of China (No. 2020SKA0110100), and, the CAS Project for Young Scientists in Basic Research (No. YSBR-062). G.L. acknowledges the support from the Natural Science Foundation of China (NSFC) (No. 12333001). We also acknowledge the IllustrisTNG collaborations for making their simulation data publicly available.